Suppose that you decided to invest $650,000 in British equity securities at a then current spot rate of $1.2745/pound. At the end of one year the spot rate is $1.3250/pound. Initially, the average price per share was GBP60, now the price is GBP57. What is your average rate for return after converting your stocks back into USD? (.....%) (Bb is not accepting percentage sign in calculated numeric questions. Thus, just type the number before percentage sign, please. If you believe the result has a negative sign, put it in front of the number)
Answer #1

investment = $650000, Current spot rate = 1.2745$ / pound

investment in Pound = investment in $ / 1.2745 = 650000 / 1.2745 = 510003.9231 pound = 510003.92 pound

No of shares brought = Investment in pound / Average share price = 510003.92 / 60 = 8500.0653 = 8500 shares (rounded as shares cannot be in decimal)

Value of investment after one year = no of shares bought x share price after one year = 8500 x 60 = 510000 pounds

Spot rate at end of 1 year = 1.3250

Value of investment after one year in $ = 510000 x 1.3250 = $675750

Average return after converting back to $ = (Value of investment after in year in $ / investment ) - 1 = (675750 / 650000) - 1 = 1.039615 - 1 = 0.039615 = 3.9615% = 3.96% (rounded to two decimal places)

Hence average return = 3.96%